OpenStudy (anonymous):

Simplify 8 to the power of negative 4 over 3

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so you need to know \[8 = 2^3\] and the question is asking you to simplify \[8^{-\frac{4}{3}} = (2^3)^{-\frac{4}{3}}\] there you go... just use the power of a power rule described in your last question.

OpenStudy (campbell_st):

oh... and you will need to know about negative powers \[x^{-a} = \frac{1}{x^a}\]

OpenStudy (anonymous):

so ?

OpenStudy (campbell_st):

I'm not answering it... I'm giving you advice on how you can answer it. Hope it helps.

OpenStudy (anonymous):

@campbell_st im in dire need of your assitance dont leave me stranded

OpenStudy (campbell_st):

so here is my work again \[8 = 2^3\] a refresher on power of a power \[(x^a)^b = x^{a \times b}\] so you have \[8^{-\frac{4}{3}} = (2^3)^{-\frac{4}{3}}\] now hopefully you can simplify it.

OpenStudy (campbell_st):

don't forget to scroll up... and look at negative powers..